# Enhancing Reasoning for Diffusion LLMs via Distribution Matching Policy Optimization

### v1

@@ B.2 Insights for Weight Baselines: Approximate Variance Reduction
Combined with ˜9, we can see that subtracting $\frac{p_{\bar{\theta}}({\bm{o}}|{\bm{q}};{\bm{\sigma}})}{p_{v}({\bm{o}}|{\bm{q}};{\bm{\sigma}})}$ from the weight does not change the gradient of the CE loss, i.e.,
@@
Theoretically, there is an optimal choice of $\lambda$ that minimizes the variance. The natural choice of $\lambda=1$ means implicitly matching the probability $p_{\theta}({\bm{o}}|{\bm{q}};{\bm{\sigma}})$ to fit $p_{*}({\bm{o}}|{\bm{q}};{\bm{\sigma}})$, which corresponds to our model weight baseline ˜15. When the frequency for sampling buffer $F$ is small, we can assume $p_{\theta}({\bm{o}}|{\bm{q}};{\bm{\sigma}})$ does not deviate too much from $p_{v}({\bm{o}}|{\bm{q}};{\bm{\sigma}})$, thus this ratio should be close to $1$, which corresponds to our group weight baseline ˜13. Finally, as we actually use the negative ELBO $\mathcal{L}_{\theta}({\bm{o}}|{\bm{q}})$ instead of $\log p_{\theta}({\bm{o}}|{\bm{q}})$ in computing the loss, the variance reduction only holds approximately.
B.3 Proofs for the Weighted Direct Discriminative Optimization Objective
@@ C.2 Training Hyperparameters and Evaluation
We use a maximum generation length $256$ tokens, a batch size of $8$ per GPU, and gradient accumulation steps of $2$, and $16$ generated rollouts per prompt. We optimized the model using the AdamW optimizer (Loshchilov & Hutter, 2019) with parameters $\beta_{1}=0.9,\beta_{2}=0.99$, weight decay of $0.1$, learning rate of $3\times 10^{-6}$, and gradient clipping at $0.2$. For each clean sequence, we sampled $4$ partially masked tokens for computing WDCE/WDDO loss. For rollouts generation during training, we use a semi-autoregressive random order sampler with a block size of $32$ to generate diverse responses, which is the recommended practice for using LLaDA series models as is described in Nie et al. (2025b). We train $4,000$ steps (number of gradient updates) for GSM8K and MATH500, Countdown, and Sudoku, respectively.
@@
For computational efficiency, we use Flash Attention 2 (Dao, 2024) and $4$-bit quantization. All experiments on DMPO share these hyperparameters. The main result reported in Tab.˜1 used the group weight baseline defined in ˜13. The ablation study in Fig.˜5 also follows the same set of hyperparameters above, except for using different choices of weight baselines.
@@
C.3 Example Outputs of the Model after Fine-tuning
@@
Model: DMPO-LLaDA Dataset: MATH500 Generation Sequence Length: $512$ Question: The product of $3t^{2}+5t+a$ and $4t^{2}+bt-2$ is $12t^{4}+26t^{3}-8t^{2}-16t+6$. What is $a+b$? To find the values of $a$ and $b$ such that the product of $3t^{2}+5t+a$ and $4t^{2}+bt-2$ equals $12t^{4}+26t^{3}-8t^{2}-16t+6$, we need to expand the left-hand side and compare coefficients with the right-hand side. First, expand the product: $(3t^{2}+5t+a)(4t^{2}+bt-2)$ Using the distributive property (FOIL method for binomials): $3t^{2}\cdot 4t^{2}+3t^{2}\cdot bt+3t^{2}\cdot(-2)+5t\cdot 4t^{2}+5t\cdot bt+5t\cdot(-2)+a\cdot 4t^{2}+a\cdot bt+a\cdot(-2)$ Simplify each term: $12t^{4}+3bt^{3}-6t^{2}+20t^{3}+5bt^{2}-10t+4at^{2}+abt-2a$ Combine like terms: $12t^{4}+(3b+20)t^{3}+(-6+5b+4a)t^{2}+(-10+ab)t-2a$ We need this to equal: $12t^{4}+26t^{3}-8t^{2}-16t+6$ By comparing coefficients, we get the following system of equations: $3b+20=26\implies 3b=6\implies b=2$ $6+5(2)+4a=-8\implies 4+4a=-8\implies 4a=-12\implies a=-3$ Thus, a=3a=-3 and $b=2$. a+b=3+2=1a+b=-3+2=-1 </reasoning><answer>1</answer>
Model: DMPO-LLaDA Dataset: GSM8K Generation Sequence Length: $512$ Question: Katarina has 5 less cookies than Max has. Max has 12 more cookies than the Cookie Monster, and Summer has 23 more cookies than Max. If Katarina has 68 cookies, how many cookies do they have in total? To determine the total number of cookies Katarina, Max, and Summer have, we need to find out how many cookies each person has and then sum them up. 1. We know that Katarina has 68 cookies. 2. Katarina has 5 less cookies than Max. 3. Max has 12 more cookies than the Cookie Monster. 4. Summer has 23 more cookies than Max. Let’s denote the number of cookies each person has as follows: - $K$ for Katarina’s cookies - $M$ for Max’s cookies - $C$ for the Cookie Monster’s cookies - $S$ for Summer’s cookies From the problem, we have: - $K=68$ - K=M5K=M-5 - $M=C+12$ - $S=M+23$ First, we substitute the known value of $K$ into the equations: - 68=M568=M-5 - $M=C+12$ - $S=M+23$ First, solve for $M$: 68=M568=M-5 $M=68+5$ $M=73$ Next, solve for $C$: $73=C+12$ C=7312C=73-12 $C=61$ Finally, solve for $S$: 73=S2373=S-23 $S=73+23$ $S=96$ Now, we have the number of cookies each person has: - Katarina: 68 cookies - Max: 73 cookies - The Cookie Monster: 61 cookies - Summer: 96 cookies To find the total number of cookies, we sum these amounts: $68+73+61+96=298$ Thus, the total number of cookies is $\boxed{298}$. </reasoning><answer>298</answer>

### v2

@@ B.3 Insights for Weight Baselines: Approximate Variance Reduction
Combined with 9, we can see that subtracting $\frac{p_{\bar{\theta}}({\bm{o}}|{\bm{q}};{\bm{\sigma}})}{p_{v}({\bm{o}}|{\bm{q}};{\bm{\sigma}})}$ from the weight does not change the gradient of the CE loss, i.e.,
@@
Theoretically, there is an optimal choice of $\lambda$ that minimizes the variance. The natural choice of $\lambda=1$ means implicitly matching the probability $p_{\theta}({\bm{o}}|{\bm{q}};{\bm{\sigma}})$ to fit $p_{*}({\bm{o}}|{\bm{q}};{\bm{\sigma}})$, which corresponds to our model weight baseline 15. When the frequency for sampling buffer $F$ is small, we can assume $p_{\theta}({\bm{o}}|{\bm{q}};{\bm{\sigma}})$ does not deviate too much from $p_{v}({\bm{o}}|{\bm{q}};{\bm{\sigma}})$, thus this ratio should be close to $1$, which corresponds to our group weight baseline 13. Finally, as we actually use the negative ELBO $\mathcal{L}_{\theta}({\bm{o}}|{\bm{q}})$ instead of $\log p_{\theta}({\bm{o}}|{\bm{q}})$ in computing the loss, the variance reduction only holds approximately.
B.4 Proofs for the Weighted Direct Discriminative Optimization Objective
@@ C.2 Training Hyperparameters and Evaluation
We use a maximum generation length $256$ tokens, a batch size of $8$ per GPU, and gradient accumulation steps of $2$, and $16$ generated rollouts per prompt. We optimized the model using the AdamW optimizer (Loshchilov and Hutter, 2019) with parameters $\beta_{1}=0.9,\beta_{2}=0.99$, weight decay of $0.1$, learning rate of $3\times 10^{-6}$, and gradient clipping at $0.2$. For each clean sequence, we sampled $4$ partially masked tokens to compute the WDCE/WDDO loss. For rollouts generation during training, we use a semi-autoregressive random order sampler (with temperature 0) and fast-dLLM (with temperature 0.2) with a block size of $32$ to generate diverse responses, which is the recommended practice for using LLaDA series models as is described in Nie et al. (2025b). We train $4,000$ steps (number of gradient updates) for GSM8K and MATH500, Countdown, and Sudoku, respectively.
In Tab. 1, DMPO-LLaDA uses the group weight baseline 13 on GSM8K, MATH500, and Sudoku, and the individual weight baseline 14 on Countdown. DMPO-LLaDA-SFT and DMPO-LLaDA-1.5 adopt the individual weight baseline 14 in all cases.
@@
For computational efficiency, we use Flash Attention 2 (Dao, 2024) and $4$-bit quantization. All experiments on DMPO share these hyperparameters. The main result reported in Tab. 1 used the group weight baseline defined in 13. The ablation study in Fig. 4 also follows the same set of hyperparameters above, except for using different choices of weight baselines.
@@
C.3 Further Experimental Results
Figure 6: Reward dynamics during training. DMPO consistently produces higher rewards than d1.
DMPO consistently achieves higher rewards
In Fig. 6, we present the reward dynamics of DMPO across training steps and compare with that of d1. DMPO consistently achieves higher reward values after an initial warm-up phase and ultimately discovers responses with higher rewards than d1, possibly because it continuously explores the reward distribution landscape throughout training. In the first $1,000$ steps, DMPO often produces lower reward values than d1, potentially due to the lack of an SFT phase before RL scaling. Moreover, we observe that the performance of DMPO does not saturate after $4,000$ gradient steps, suggesting its greater potential than GRPO-type algorithms.
Ablation studies on the hyperparameter dependence
We provide an ablation study on two of the main hyperparameters in Alg. 1, namely the number of rollouts $N$ and the frequency for sampling buffer $F$, in Figs. 7 and 8, respectively. For each run shown in Fig. 7, we train for $6$ hours using $8$ NVIDIA H200 GPUs. For each run shown in Fig. 8, we train for $8$ hours with $8$ NVIDIA H200 GPUs. We only vary the resampling buffer frequency $F$ and the number of rollouts sampled per prompt $N$, while fixing other hyperparameters, such as the total effective batch size, to maintain a fair comparison.
For the number of rollouts per question $N$, we observe that a larger number of $N$ does not necessarily lead to longer training time, even with the same number of steps, due to the parallelism of the generation process, since we kept the total batch size fixed while varying the hyperparameter $N$. The algorithm is robust across various values of $N$ ranging from $4$ to $32$ thanks to the mechanism for inserting negative gradients.
For the buffer sampling frequency $F$, we observe that it significantly affects training speed. The figure clearly demonstrates the advantage of DMPO due to its off-policy nature, whereas a purely on-policy realization of WDCE loss (with $F=1$) is not only extremely slow but also does not show a significant boost in per-step reward gains. The figure also underscores the unique benefit of WDCE being a forward loss: given the generated rollouts and their weights, one can train using the simple forward process via random masking. Our algorithm is robust to choices of $F$ up to $24$, whereas an even larger $F$ may cause slight instability later in training when the reward is high.
Figure 7: Ablation study of the number of rollouts per prompt NN on Countdown dataset under the same training time and compute. The performance is robust to this hyperparameter.
Figure 8: Ablation study of the resampling frequency $F$ on Countdown dataset. A larger $F$ is generally more time-efficient though may cause instability when the reward is high.
Visualizing rollout entropy of DMPO
In Fig. 9, we compare the reward and entropy of the generated rollouts during training for both the relative-entropy-based (diffu-GRPO) RL algorithm and the cross-entropy-based (DMPO) RL algorithm. Here, in both experiments, we fix $N=16$ and $F=8$ and evaluate the entropy of generated samples every $10$ generations. The evaluation of entropy is as follows: we use random-order autoregressive generation with block length $32$, and at the $d$-th step of unmasking (where $d$ ranges from 11 to $D=|{\bm{o}}|$), we compute the entropy of the predicted logits at the $d$-th position, and take average of all the $D$ entropy values as the final value of sequential entropy. From the figure, the trend of consistently higher sample entropy for WDCE loss than for diffu-GRPO agrees with our expectation that cross-entropy-based methods are less prone to mode-seeking and maintain a higher level of diversity throughout training.
Figure 9: Reward (left) and entropy (right) of generated rollouts during training for both diffu-GRPO (Zhao et al., 2025a) and DMPO.
C.4 Example Outputs of the Model after Fine-tuning
@@
[…]